# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 6.5 Evaluation Methodologies and Benchmarking for Merged Model Assessment
The diverse application scenarios examined in this section demonstrate that model merging has matured from a theoretical curiosity into a practical paradigm with tangible impact across capability augmentation, safety alignment, federated learning, and domain specialization, while the evaluation methodologies discussed provide essential frameworks for rigorously assessing merged model quality. These real-world deployments and assessment practices naturally raise questions about the supporting infrastructure, persistent limitations, and emerging opportunities that will shape the field’s evolution. Beyond these specific application scenarios, the continued advancement of model merging relies heavily on the broader ecosystem of tools, benchmarks, and community resources. Table 8 summarizes these application scenarios. The following section examines this infrastructure, alongside a synthesis of open challenges and future research directions that will determine the trajectory of model merging as a foundational paradigm. Table 5 provides a cross-dimensional comparison of all methods.
@@ 7.1 Summary of Contributions
With this foundation in place, we now turn to synthesizing the broader state of model merging research. This survey has presented a comprehensive examination of model merging in the era of large language models, systematically organizing the rapidly expanding landscape of techniques, theoretical foundations, and practical applications. Our central contribution is the unified taxonomic framework that categorizes merging methodologies along multiple dimensions (Table 5 summarizes the characteristics of each category): weight-space averaging and geometric interpolation methods that treat parameters as points in Euclidean or manifold spaces; task vector arithmetic and sparsification-enhanced approaches that conceptualize fine-tuning as composable vector displacements; and structured information-guided methods that preserve architectural distinctions or leverage runtime statistics for principled combination. Beyond methodological categorization, we have synthesizedd the theoretical underpinnings that explain merging success, including loss landscape geometry, linear mode connectivity (Singh and Jaggi, 2020), and weight space symmetries, providing researchers with the conceptual tools necessary to understand when and why merging succeeds. We have further connected these methods to diverse application scenarios spanning capability augmentation, alignment and safety, and efficiency-driven distributed learning contexts. This comprehensive mapping of the methodological and theoretical landscape provides the vantage point from which to identify the key trends that have characterized the field’s technical evolution, as depicted in Figure 4.
@@ 7.2 Technical Evolution Trends
Looking across the methodological landscape, several notable trends characterize the technical evolution of model merging. First, the field has progressed from simple uniform averaging toward increasingly sophisticated interference-aware methods that explicitly address sign conflicts and parameter redundancy through sparsification and consensus mechanisms. Second, there has been a marked shift from static, manually-specified merging configurations toward automated and evolutionary approaches that discover optimal recipes through search and optimization. Third, the community has increasingly recognized the importance of representation-level information, moving beyond purely weight-space operations to incorporate activation statistics and functional correspondence into merging decisions. Fourth, the scale of merging applications has expanded dramatically, from combining a handful of fine-tuned variants to orchestrating complex multi-model compositions that integrate dozens of specialized experts. Finally, the tight coupling between theoretical understanding and algorithmic innovation has accelerated, with insights from mode connectivity theory directly informing practical method design. While these trends collectively demonstrate remarkable progress toward principled model combination, they also illuminate the boundaries of current understanding and point toward substantial challenges that continue to limit the field’s advancement.
@@ 7.3 Open Challenges and Limitations
Several open challenges remain. First, a rigorous theoretical explanation for why large pretrained models exhibit such favorable mergeability is still lacking; current understanding relies on empirical observations rather than formal guarantees (He et al., 2025; Nobari et al., 2025). Second, scalability poses a significant barrier: as models grow to hundreds of billions of parameters, alignment and conflict resolution costs scale super-linearly. Third, standardized evaluation protocols remain absent—the field lacks consensus benchmarks that measure both individual task retention and cross-task transfer after merging. Fourth, the absence of established best practices for selecting merge configurations (which models to merge, at what granularity, with which algorithm) hinders practitioners’ ability to make principled decisions.
@@ 7.4 Future Research Directions
The open challenges identified above naturally point toward several promising research frontiers that could significantly advance the field of model merging.
@@
Among the most immediately impactful of these emerging research directions is developing systems that can automatically predict merging outcomes and recommend optimal configurations without exhaustive experimentation. Current approaches require substantial trial-and-error to identify effective merge recipes, representing a significant barrier to practical adoption. Future work should focus on learning predictive models that estimate merged model performance from properties of constituent models, such as gradient alignment, representation similarity, and task relatedness, enabling rapid configuration selection. The formal objective can be expressed as:
$\mathcal{C}(\theta_{1},\ldots,\theta_{k})=f_{\phi}(\text{Sim}(\theta_{i},\theta_{j}),\text{Task}(\theta_{i}),\text{Arch}(\theta_{i}))\rightarrow\mathbb{E}[\text{Perf}(\text{Merge}(\theta_{1},\ldots,\theta_{k}))]$ (11)
where $f_{\phi}$ is a learned compatibility function parameterized by $\phi$. Such systems could substantially reduce the experimental overhead of merging while improving accessibility for practitioners without deep expertise in merging methodology.
@@
Extending merging capabilities beyond architecturally identical models represents a transformative research opportunity. Current constraints requiring shared architecture and initialization exclude vast repositories of independently developed models from participation in merging workflows. Key research directions include:
  • Representation-level translation mechanisms: Learning mappings between feature spaces of different architectures
  • Architecture-agnostic knowledge distillation (Hinton et al., 2015): Transferring capabilities without requiring parameter correspondence
  • Learned correspondence mappings: Establishing soft alignments between functionally equivalent components across model families
  • Hybrid merging strategies: Combining parameter-level and representation-level operations for heterogeneous models
@@
Real-world deployment scenarios increasingly require models that can continuously incorporate new capabilities without full retraining. Dynamic merging systems that can incrementally integrate new expert models, adapt merging coefficients based on runtime performance feedback, and handle concept drift in constituent model capabilities represent an underexplored frontier. The mathematical formulation for continual merging extends the static case:
$\theta_{\text{merged}}^{(t+1)}=\text{Update}(\theta_{\text{merged}}^{(t)},\theta_{\text{new}},\mathcal{D}_{\text{val}})$ (12)
@@
Establishing rigorous theoretical foundations for model merging remains a critical priority. Key open questions include:
@@
Recent work has begun exploring merging in alternative representational spaces beyond the standard weight domain. FREE-Merging (Zheng and Wang, 2025) proposes performing model merging in the Fourier transform domain, decomposing task vectors into frequency components and applying lightweight expert modules to selectively recombine them, achieving state-of-the-art results on multi-task benchmarks (ICCV 2025). Concurrently, Fu et al. (2025) introduce model merging as a mechanism for knowledge editing, demonstrating that targeted task vector manipulations can insert, update, or erase factual knowledge in LLMs without retraining. These directions suggest that the conceptual toolkit for model merging extends well beyond simple parameter averaging.
@@
As merged models see increasing deployment in safety-critical applications, developing merging frameworks that provide formal guarantees about alignment preservation becomes essential. Future work should investigate:
@@ 7.5 Outlook
Model merging has evolved from simple checkpoint averaging into a sophisticated paradigm for composing specialized capabilities into unified systems. The proliferation of openly available fine-tuned LLMs has created unprecedented opportunities for leveraging merging techniques to democratize access to multi-capability AI. The community has developed.[truncated]
Looking forward, we anticipate that model merging will become an increasingly central component of the machine learning pipeline, with several key trends shaping the field’s trajectory:
@@
    Integration with Automated ML: Merging configuration discovery will become increasingly automated through neural architecture search and hyperparameter optimization techniques, reducing the expertise barrier for practitioners.
@@
    Specialized Tooling and Infrastructure: The emergence of dedicated merging platforms and standardized interfaces will accelerate adoption and enable more sophisticated merging workflows.
@@
    Theoretical Maturation: As empirical understanding deepens, we expect the development of rigorous theoretical frameworks that provide guarantees about merged model behavior and enable principled design decisions.
@@
    Multimodal and Cross-Domain Merging: Techniques for combining models across modalities (text, vision, audio) and domains will unlock new capabilities for building comprehensive AI assistants.
@@
    Safety and Alignment Integration: Merging frameworks will increasingly incorporate safety constraints as first-class citizens, ensuring that capability enhancement does not compromise alignment properties.
In conclusion, model merging represents a fundamental change in how we think about building and deploying large language models. Rather than training monolithic models from scratch for each application, the field is moving toward a compositional approach where specialized capabilities can be flexibly combined to meet diverse requirements. This survey has aimed to provide researchers and practitioners with a comprehensive foundation for understanding the current state of the art and contributing to the continued advancement of this transformative technology.
@@ 8 Conclusion
This survey has organized the rapidly expanding model merging landscape through a unified FUSE taxonomy—Foundations, Unification Strategies, Scenarios, and Ecosystem—covering the theoretical, algorithmic, practical, and infrastructural dimensions of the field.
On the algorithmic front, we traced the evolution from simple weight averaging and Model Soups (Wortsman et al., 2022a) through task-vector arithmetic (Ilharco et al., 2023) and interference-aware methods such as TIES-Merging (Yadav et al., 2023), DARE (Yu et al., 2023), and DELLA (Deep et al., 2024), to adaptive strategies like AdaMerging (Yang et al., 2024d) and evolutionary optimization (Akiba et al., 2024). As summarized in Table 5, each family of methods navigates a fundamental trade-off between the simplicity of data-free merging and the accuracy of data-dependent or search-based strategies.
On the theoretical side, we synthesized results from loss landscape geometry, linear mode connectivity (Singh and Jaggi, 2020), and permutation symmetry that collectively explain when and why merging succeeds. These insights are not merely explanatory: they directly inform practical algorithm design, as witnessed by the progression from naïve averaging to curvature-aware and sign-resolved merging.
On the applications front, we surveyed how merging enables multi-task generalization, multilingual transfer, safety alignment, and federated learning—all at a fraction of the cost of full retraining. However, we also highlighted dual-use risks, evaluation gaps, and the absence of standardized benchmarks (He et al., 2025) as pressing concerns.
Looking ahead, we identified scalability to frontier-scale models, cross-architecture and cross-modality merging, dynamic continual merging, and safety-preserving merge guarantees as the most impactful open problems. As the number of publicly available fine-tuned models continues to grow, principled merging techniques will become increasingly central to the AI practitioner’s toolkit. We hope this survey helps researchers and practitioners navigate this rapidly evolving field.

### 30 Mar 2026 v2

@@ 6.5 Evaluation Methodologies and Benchmarking for Merged Model Assessment
These real-world deployments and assessment practices naturally raise questions about the supporting infrastructure, persistent limitations, and new opportunities that will shape the field’s evolution. Table 8 summarizes these application scenarios. Building on this foundation, the final section synthesizes open challenges, future research directions, and the ecosystem infrastructure that will determine the trajectory of model merging. Table 5 provides a cross-dimensional comparison of all methods.
@@ 7.1 Summary of Contributions
This survey has presented a structured examination of model merging in the era of large language models, organizing the rapidly expanding body of techniques, theoretical foundations, and practical applications through the FUSE taxonomy. Our central contribution is the unified taxonomic framework that categorizes merging methodologies along multiple dimensions (Table 5 summarizes each category); weight-space averaging and geometric interpolation methods that treat parameters as points in Euclidean or manifold spaces; task vector arithmetic and sparsification-enhanced approaches that conceptualize fine-tuning as composable vector displacements; and structured information-guided methods that preserve architectural distinctions or exploit runtime statistics for informed combination. Beyond methodological categorization, we have synthesized the theoretical underpinnings explaining merging success, including loss surface geometry, linear mode connectivity (Frankle et al., 2020), and weight space symmetries, and connected these methods to diverse application scenarios spanning capability augmentation, alignment and safety, and efficiency-driven distributed learning. This mapping of the methodological and theoretical terrain provides the vantage point from which to identify the key trends in the field’s technical evolution, as depicted in Figure 4.
@@ 7.2 Technical Evolution Trends
Five interconnected trends characterize the methodological evolution of model merging. First, the field has progressed from simple uniform averaging toward interference-aware methods that explicitly resolve sign conflicts and parameter redundancy through sparsification and consensus mechanisms. Second, there has been a shift from static, manually-specified configurations toward automated and evolutionary approaches that discover optimal recipes through search. Third, researchers have increasingly incorporated representation-level information, moving beyond purely weight-space operations to use activation statistics and functional correspondence in merging decisions. Fourth, the scale of applications has expanded from combining a handful of fine-tuned variants to orchestrating multi-model compositions integrating dozens of specialized experts. Fifth, the coupling between theoretical understanding and algorithmic innovation has tightened, with insights from mode connectivity theory directly informing practical method design. While these trends mark substantial progress, they also illuminate the boundaries of current understanding and the open challenges that constrain further advancement.
@@ 7.3 Open Challenges and Limitations
Several open challenges remain. First, a rigorous theoretical explanation for why large pretrained models exhibit such favorable mergeability is still lacking; current understanding relies on empirical observations rather than formal guarantees (He et al., 2025b; Nobari et al., 2025), though recent work on certifiable merging (Kim et al., 2025) shows that non-vacuous generalization bounds can be derived for merged models, representing a promising step toward formal guarantees. Second, scalability poses a major barrier, since as models grow to hundreds of billions of parameters, alignment and conflict resolution costs scale super-linearly; approaches like Navigating the Accuracy-Size Trade-Off (Dhasade et al., 2025) and cross-architecture task vector transport (Rinaldi et al., 2026) begin to tackle this but remain at early stages. Third, standardized evaluation protocols remain absent; the field lacks consensus benchmarks that measure both individual task retention and cross-task transfer after merging, though FusionBench (Tang et al., 2024c) and MergeBench (He et al., 2025b) have made notable strides. Fourth, the absence of established best practices for selecting merge configurations (which models to merge, at what granularity, with which algorithm) hinders practitioners’ ability to make principled decisions, motivating tools such as MergePipe (Wang et al., 2026c) for budget-aware parameter management.
7.4 Systematic Failure Modes of Model Merging
While the preceding sections have emphasized the successes of model merging, a complete understanding requires systematic analysis of when and why merging fails. We identify four principal failure modes, each with distinct causes, symptoms, and diagnostic indicators.
Task-level merging collapse.
Cao et al. (2026b) identify and characterize a fundamental failure mode they term merging collapse: certain combinations of task-specialist models suffer catastrophic performance degradation after merging, regardless of the merging method applied. Through extensive experiments, they demonstrate that representational incompatibility between tasks, measured by the divergence of hidden-state distributions across models, is strongly correlated with collapse, while parameter-space conflict metrics (e.g., sign disagreement rates) show minimal correlation. This finding challenges the conventional wisdom that parameter interference is the primary cause of merging failure, and suggests that representational diagnostics should take precedence in mergeability assessment. They further provide theoretical bounds via rate-distortion theory, establishing fundamental limits on task mergeability that hold regardless of methodology.
Parameter interference and sign conflicts.
Even when tasks are representationally compatible, parameter-level conflicts can degrade performance. TIES-Merging (Yadav et al., 2023) demonstrated that when task vectors for different tasks have opposing signs at the same parameter positions, naïve averaging can produce destructive cancellation that nullifies the contributions of both tasks. TSV-Merge (Gargiulo et al., 2025) further shows that task interference operates at the level of singular value structure: irrelevant singular vectors from one task can corrupt the subspace of another, and reducing this cross-task interference via whitening transformations yields measurable improvements. The practical implication is that sign conflict rates and subspace overlap metrics should be monitored as early indicators of potential degradation.
Mergeability prediction failures.
Zhou et al. (2026) show that mergeability is not an intrinsic property of model pairs but depends fundamentally on both the merging method and the partner tasks. Their analysis reveals that the properties predicting successful merging vary substantially across methods (only 46.7% metric overlap and 55.3% sign agreement between methods), revealing method-specific “fingerprints” for compatibility. This means that a pair of models that merges well with one method may fail with another, and practitioners cannot rely on a single diagnostic to predict outcomes. However, subspace overlap and gradient alignment metrics consistently emerge as foundational, method-agnostic prerequisites for compatibility.
Evaluation-masked failures.
Tam et al. (2024b) provide a cautionary finding: under realistic evaluation conditions that test compositional generalization (i.e., requiring the merged model to handle combinations of capabilities not explicitly present in any source model), the reported gains from many merging methods diminish substantially or fail to transfer. Standard evaluation protocols that test each capability in isolation may overstate merging effectiveness by failing to detect subtle interference between merged capabilities. This highlights the need for evaluation protocols that specifically probe cross-capability interaction effects in merged models.
Taken together, these failure modes suggest a diagnostic workflow for practitioners. First, (1) assess representational compatibility between candidate models before merging to screen for potential collapse; (2) monitor parameter-level sign conflicts and subspace overlap as secondary indicators; (3) match diagnostic metrics to the specific merging method being used, since mergeability predictors are method-dependent; and (4) evaluate merged models using compositional generalization benchmarks rather than single-capability tests alone.
@@ 7.5 Future Research Directions
The open challenges identified above naturally point toward several promising research frontiers that could substantially advance the field of model merging.
@@
Among the most immediately impactful of these emerging research directions is developing systems that can automatically predict merging outcomes and recommend optimal configurations without exhaustive experimentation. Current approaches require considerable trial-and-error to identify effective merge recipes, representing a barrier to practical adoption. SimMerge (Bolton et al., 2026) takes a step in this direction by learning to select merge operators from similarity signals, while the investigation of mergeability causes (Rahamim et al., 2026) provides foundational understanding of when and why models can be successfully combined. PSO-Merging (Zhang et al., 2025d) applies particle swarm optimization, AIMMerging (Feng et al., 2025) uses adaptive iterative merging along training trajectories, Bayesian model-merging previews (Maldonado et al., 2024) allow fast previewing of multi-task fine-tuning outcomes, and black-box model merging (Chen et al., 2025c) tackles the LLM-as-a-Service scenario with massive model repositories. Future work should focus on learning predictive models that estimate merged model performance from properties of constituent models, such as gradient alignment, representation similarity, and task relatedness, allowing rapid configuration selection. The formal objective can be expressed as:
$\mathcal{C}(\theta_{1},\ldots,\theta_{k})=f_{\phi}(\text{Sim}(\theta_{i},\theta_{j}),\text{Task}(\theta_{i}),\text{Arch}(\theta_{i}))\approx\mathbb{E}[\text{Perf}(\text{Merge}(\theta_{1},\ldots,\theta_{k}))],$ (11)
where $f_{\phi}$ is a learned compatibility function parameterized by $\phi$. Such systems could greatly reduce the experimental overhead of merging while improving accessibility for practitioners without deep expertise in merging methodology.
@@
Extending merging capabilities beyond architecturally identical models represents an important research opportunity. Current constraints requiring shared architecture and initialization exclude vast repositories of independently developed models from participation in merging workflows. Transport and Merge (Cui et al., 2026) takes a pioneering step by proposing optimal transport-based methods for cross-architecture merging of large language models, showing that models with different layer configurations can be meaningfully combined. AdaMMS (Du et al., 2025) tackles merging of heterogeneous multimodal large language models with unsupervised coefficient optimization, extending merging beyond text-only models. Key research directions include:
  • Representation-level translation mechanisms. Learning mappings between feature spaces of different architectures
  • Architecture-agnostic knowledge distillation (Hinton et al., 2015). Transferring capabilities without requiring parameter correspondence
  • Learned correspondence mappings. Establishing soft alignments between functionally equivalent components across model families
  • Hybrid merging strategies. Combining parameter-level and representation-level operations for heterogeneous models
@@
Real-world deployment scenarios increasingly require models that can continuously incorporate new capabilities without full retraining. Dynamic merging systems that can incrementally integrate new expert models, adapt merging coefficients based on runtime performance feedback, and handle concept drift in constituent model capabilities represent an underexplored frontier. Kleiman et al. (2025) show that model averaging provides an effective mechanism for mitigating forgetting during continual learning, establishing “Soup to go” as a practical approach for maintaining previously acquired knowledge while integrating new capabilities. Scalable Model Merging (Xu et al., 2025b) tackles the scalability challenge through progressive layer-wise distillation, supporting efficient merging of increasingly large models. RAIN-Merging (Huang et al., 2026b) provides a gradient-free method specifically designed for enhancing instruction following in large reasoning models while preserving thinking format, highlighting the emerging importance of reasoning-aware merging. The mathematical formulation for continual merging extends the static case:
$\theta_{\text{merged}}^{(t+1)}=\text{Update}(\theta_{\text{merged}}^{(t)},\theta_{\text{new}},\mathcal{D}_{\text{val}}),$ (12)
@@
Establishing rigorous theoretical foundations for model merging remains a critical priority. Recent progress includes certifiable merging bounds (Kim et al., 2025) that provide non-vacuous generalization guarantees, theoretical analysis of task vector effectiveness in nonlinear Transformers (Li et al., 2025b), and empirical-theoretical studies of task-level model-merging collapse (Cao et al., 2026b). On the fairness front, Naganuma et al. (2025) examine the role of task vectors in ensuring equitable performance across subgroups. Key open questions include:
@@
Recent work has begun exploring merging in alternative representational spaces beyond the standard weight domain. FREE-Merging (Zheng and Wang, 2025) proposes performing model merging in the Fourier transform domain, decomposing task vectors into frequency components and applying lightweight expert modules to selectively recombine them, achieving top results on multi-task benchmarks (ICCV 2025). Concurrently, Fu et al. (2025b) introduce model merging as a mechanism for knowledge editing, combining supervised fine-tuning on new knowledge with subsequent model merging to preserve both updated facts and general capabilities without architectural changes. Weighted-Reward Preference Optimization (Yang et al., 2025d) explores implicit model fusion through reward-weighted preference signals. Bridging Training and Merging (Moayedikia and Troncoso, 2026) proposes momentum-aware optimization that unifies the training and merging phases. Model Merging via Multi-Teacher Knowledge Distillation (Dalili and Mahdavi, 2025) bridges the gap between parameter-level merging and output-level knowledge transfer. A Unified View of Delta Parameter Editing (Tang et al., 2024f) provides a unified framework for understanding post-training parameter modifications. These directions suggest that the conceptual toolkit for model merging extends well beyond simple parameter averaging.
@@
As merged models see increasing deployment in safety-critical applications, developing merging frameworks that provide formal guarantees about alignment preservation becomes essential. Among Us (Yang et al., 2026d) confronts the critical challenge of measuring and mitigating malicious contributions in model collaboration systems, where adversarial participants may attempt to inject harmful behaviors through the merging process. Future work should investigate:
@@ 7.6 Outlook
Model merging has evolved from simple checkpoint averaging into a versatile methodology for composing specialized capabilities into unified systems. The proliferation of openly available fine-tuned LLMs creates rich opportunities for merging techniques to democratize access to multi-capability AI. We anticipate five key trends shaping the field’s trajectory. First,
@@
    Integration with Automated ML. Merging configuration discovery will become increasingly automated through neural architecture search and hyperparameter optimization techniques (Lu et al., 2026; Su and Geiping, 2025), reducing the expertise barrier for practitioners.
@@
    Specialized Tooling and Infrastructure. Dedicated merging platforms such as MergeKit (Goddard et al., 2024), Mergenetic (Minut et al., 2025), and budget-aware management systems (Wang et al., 2026c) will accelerate adoption and enable more advanced merging workflows.
@@
    Theoretical Maturation. As empirical understanding deepens, we expect rigorous theoretical frameworks that provide guarantees about merged model behavior and enable principled design decisions.
@@
    Multimodal and Cross-Domain Merging. Techniques for combining models across modalities (text, vision, audio) and domains will unlock new capabilities for building general-purpose AI assistants. Recent works such as MergeVLA (Fu et al., 2025a) for vision-language-action agents, Bring Reason to Vision (Chen et al., 2025d) for perception-reasoning integration, and training-free multimodal fusion (Chen et al., 2024b) for enhancing perception capabilities show the growing feasibility of cross-modal merging.
@@
    Safety and Alignment Integration. Merging frameworks will increasingly treat safety constraints as first-class citizens, ensuring that capability enhancement does not compromise alignment properties.
@@ 8 Conclusion
This survey has organized the rapidly expanding model merging field through the FUSE taxonomy (Foundations, Unification Strategies, Scenarios, and Ecosystem), providing a unified lens on the theoretical, algorithmic, practical, and infrastructural dimensions of the field.
On the algorithmic front, we traced an arc from simple weight averaging and Model Soups (Wortsman et al., 2022a) through task-vector arithmetic (Ilharco et al., 2023) and interference-aware methods such as TIES-Merging (Yadav et al., 2023) and DARE (Yu et al., 2024), to adaptive strategies like AdaMerging (Yang et al., 2024d) and evolutionary optimization (Akiba et al., 2025). Recent spectral (Lee et al., 2025c), activation-guided (Yao et al., 2025), and curvature-informed (Mahdavinia et al., 2025) techniques continue to push the frontier. As summarized in Table 5, each family navigates the fundamental tension between data-free simplicity and data-dependent precision.
On the theoretical side, we synthesized results from loss surface geometry, linear mode connectivity (Frankle et al., 2020), and permutation symmetry that collectively explain when and why merging succeeds. These insights feed directly into algorithm design. The progression from naïve averaging to curvature-aware and sign-resolved merging is, at its core, a story of theory informing practice.
On the applications front, we surveyed how merging supports multi-task generalization, multilingual transfer, safety alignment, and federated learning, all at a fraction of the cost of full retraining, while also highlighting dual-use risks, evaluation gaps, and the absence of standardized benchmarks (He et al., 2025b) as pressing concerns. We further provided a systematic analysis of failure modes (Section 7.4), cataloging when and why merging fails, from task-level collapse and sign conflicts to evaluation-masked degradation, and proposed a diagnostic workflow for practitioners.
Looking ahead, the most impactful open problems include scalability to frontier-scale models, cross-architecture merging (Cui et al., 2026), dynamic continual merging (Kleiman et al., 2025), and provably safety-preserving merge guarantees (Yuan et al., 2025). As publicly available fine-tuned models continue to proliferate, principled merging techniques will become an increasingly central part of the AI practitioner’s toolkit. We hope this survey equips both researchers and practitioners to navigate and advance this rapidly evolving field.